# Extending Chevalley’s Theorem: A Topological Characterization of Constructibility and Its Generalization Beyond Noetherian Spaces

### v1

@@ 4.3. Locally Finite Type is Weak Good
Recall that a topological space is Jacobson iff every non‑empty locally closed subset contains a closed point.
Let $U\subseteq X$ be a non‑empty locally closed subset. By Proposition 4.5, $f$ is weak good, so there exists a non‑empty locally closed subset $V\subseteq Y$ with $V\subseteq f(U)$. Since $Y$ is Jacobson, $V$ contains a closed point $y$.
The fibre $f^{-1}(y)\cap U$ is non‑empty because $y\in f(U)$. Moreover, $f^{-1}(y)$ is locally closed in $X$ (it is the inverse image of the closed point $y$), hence $f^{-1}(y)\cap U$ is locally closed in $X$. As $f$ is locally of finite type, the fibre $f^{-1}(y)$ is a scheme of finite type over the residue field $\kappa(y)$. Schemes of finite type over a field are Jacobson; therefore the non‑empty locally closed subset $f^{-1}(y)\cap U$ contains a point $x$ that is closed in $f^{-1}(y)$.
@@
5. Outlook
5.1. Formal Completions
For schemes, the good property behaves well under formal completion along closed subschemes, providing a bridge to formal geometry. More precisely:

### v2

@@ 4.4 Jacobson Ascent
Let $U\subseteq X$ be a non-empty locally closed subset. By Theorem 4.1 $f$ is good, hence weak good. Therefore there exists a non-empty locally closed subset $V\subseteq Y$ with $V\subseteq f(U)$. Since $Y$ is Jacobson, $V$ contains a closed point $y$.
The fibre $f^{-1}(y)\cap U$ is non-empty because $y\in f(U)$. Moreover, $f^{-1}(y)$ is locally closed in $X$ (it is the inverse image of the closed point $y$), hence $f^{-1}(y)\cap U$ is locally closed in $X$. As $f$ is locally of finite type, the fibre $f^{-1}(y)$ is a scheme of finite type over the residue field $\kappa(y)$. Schemes of finite type over a field are Jacobson; therefore the non-empty locally closed subset $f^{-1}(y)\cap U$ contains a point $x$ that is closed in $f^{-1}(y)$.
@@
5 Outlook
5.1 Formal Completions
@@ 5.2 A General Pattern
5.3 Extension to Non-Archimedean Geometry
Because the definition of a good map is purely topological and the key local-on-source/target properties only require sobriety, the theory developed here can be transferred directly to adic spaces and perfectoid spaces, which are sober topological spaces. A natural next step is to investigate whether important classes of morphisms in rigid-analytic or perfectoid geometry satisfy the good property. For instance:
  • Are finite-type morphisms of adic spaces good?
  • Are pro-étale morphisms or diamonds good?
  • Does an analogue of Theorem 4.1 hold in the category of perfectoid spaces?
Positive answers would yield topological generalizations of Chevalley-type results in non-Archimedean geometry. Moreover, the stability of goodness under formal completion (Proposition 5.1) suggests a natural bridge between good maps and formal geometry in the sense of Scholze-Weinstein.